Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates:

Machine-hours required to support estimated production 157,000
Fixed manufacturing overhead cost $ 658,000
Variable manufacturing overhead cost per machine-hour $ 4.60

Required:

1. Compute the plantwide predetermined overhead rate.

2. During the year, Job 400 was started and completed. The following information was available with respect to this job:

Direct materials $ 350
Direct labor cost $ 240
Machine-hours used 30

Compute the total manufacturing cost assigned to Job 400.

3. If Job 400 includes 50 units, what is the unit product cost for this job?

4. If Moody uses a markup percentage of 110% of its total manufacturing cost, then what selling price per unit would it have established for Job 400?
Answer #1

a. Predetermine Overhead rate =(658000/157000)+4.60 =4.19+4.60 =8.80(app) 350 b. Total manufacturing cost Direct material $ D